Ask
your own question, for FREE!
Mathematics
11 Online
Michelle draws a card from a standard deck of 52 cards. She replaces the card and draws a second card. What is the probability that she draws a red card and then a black card?
Still Need Help?
Join the QuestionCove community and study together with friends!
First probability = 26 red cards/52 cards = 1/2 Second probability = 26 black cards/51 cards Total probability: 1/2*26/51 = 13/51
that answer choice isn't on the test
Opps, since she replaces the cards, it's 26 black cards/ 52 CARDS, which is 1/2*1/2 = 1/4 Sorry about that... too many questions
thats okay .. and thanks alot !
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
9 hours ago
0 Replies
0 Medals
4 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
1 day ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal